Constants Part A Two people, one of mass 88 kg and the other of mass 65 kg, sit in a rowboat of mass 61 kg. With the boat initially at rest, the two people, who have been sitting at opposite ends of the boat 3.3 m apart from each other, now exchange seats. How far will the boat move? Express your answer to two significant figures and include the appropriate units. dValue Units Submit t Ans Part B In what direction will the boat move? O toward the initial position of the 65-kg person toward the initial position of the 88-kg person Submit Provide Feedback
Answer #1

Let the person with mass 88kg be at x=0.

Center of mass of the whole system is

x=\frac{88\times 0+61\times \frac{3.3}{2}+65\times 3.3}{88+61+65}\\ \implies x=1.47cm......(1)

After interchange

x'=\frac{88\times 3.3+61\times \frac{3.3}{2}+65\times 0}{88+61+65}\\ \implies x'=1.83cm

As no external force acts on the system so the boat moves by 1.83m-1.47m=0.36m to keep center of mass unchanged.

b) Towards the initial position of the person with mass 88kg